Grouper Corp. purchased a $100,000 face-value bond of Myers Corp. on August 31, 2019, for $108,980 plus accrued interest. The bond pays interest annually each November 1 at a rate of 9%. On November 1, 2019, Grouper Corp. received the annual interest. On December 31, 2019, Grouper’s year end, the fair value for these bonds was 108.1. Grouper sold the bond on January 15, 2020, for $107,800 plus accrued interest. Assume Grouper Corp. follows IFRS.

Prepare the journal entries to record the purchase of the bond, the receipt of interest, any adjustments required at year end, and the subsequent sale of the bond.

How many months was the bond held for by Grouper Corp. in 2019? Based on this, how much of the income reported on this bond should be for interest received?

How much return did Grouper Corp. earn while the bond was held?
Answer #1

Prepare the journal entries to record the purchase of the bond, the receipt of interest, any adjustments required at year end, and the subsequent sale of the bond.

date account debit credit
August 31, 2019 investment in bonds 100000
accrued interest receivable 8980
cash 108980
November 1st, 2019 cash 2250
interest revenue 2250
Dec 31, 2019 unrealized earning 8100
investment in bonds 8100
january 15 cash 107800
bonds payable 100000
interest payable 7800

How many months was the bond held for by Grouper Corp. in 2019?

It was held from august 31 to January 15th, (4 months, 15 days).

Based on this, how much of the income reported on this bond should be for interest received?

$3,375 this is the amount of interest received during the period of time the bond was held.

  • Annual interest rate 9%.
  • 100,000*9%=9000 annual interest.
  • 9000/12=750 monthly interest
  • 750*4.5months=$3,375.

How much return did Grouper Corp. earn while the bond was held?

Based on the nominal yield, we can say that the return rate was 0.09%, obtained by dividing the annual interest by the face value. Simply put, the return is equal to the interest received.